### v1

# SGFormer: Spherical Geometry Transformer for $360^{\circ}$ Depth Estimation

Abstract.
Panoramic distortion poses a significant challenge in $360^{\circ}$ depth estimation, particularly pronounced at the north and south poles. Existing methods either adopt a bi-projection fusion strategy to remove distortions or model long-range dependencies to capture global structures, which can result in either unclear structure or insufficient local perception. In this paper, we propose a spherical geometry transformer, named SGFormer, to address the above issues, with an innovative step to integrate spherical geometric priors into vision transformers. To this end, we retarget the transformer decoder to a spherical prior decoder (termed SPDecoder), which endeavors to uphold the integrity of spherical structures during decoding. Concretely, we leverage bipolar re-projection, circular rotation, and curve local embedding to preserve the spherical characteristics of equidistortion, continuity, and surface distance, respectively. Furthermore, we present a query-based global conditional position embedding to compensate for spatial structure at varying resolutions. It not only boosts the global perception of spatial position but also sharpens the depth structure across different patches. Finally, we conduct extensive experiments on popular benchmarks, demonstrating our superiority over state-of-the-art solutions.
Depth Estimation, Panoramic Distortion, Spherical Priors, Position Embedding
1. Introduction
Figure 1. (a) Bi-projection fusion strategy. Bifuse (Wang et al., 2020) combines feature maps from equirectangular and cubemap projections. (b) Long-range dependency strategy. EGFormer (Yun et al., 2023) adapts vertical and horizontal shaped local windows. (c) Our SGFormer presents a query-based global conditional position embedding(GCPE) and integrates spherical geometric priors, significantly improving the performance of panoramic depth estimation.
The panoramas represent a whole scene as an equirectangular map, delivering a $360^{\circ}$ Field of View (FoV) of the surroundings. Hence, estimating depth from a single panoramic image can offer complete depth of the scene, with a holistic understanding of the surroundings. However, the equidistant unfolding characteristic (also known as panoramic distortion) of panoramas poses a new challenge for $360^{\circ}$ depth estimation. Recently, researchers (Wang et al., 2020; Ai et al., 2023) have preferred to estimate accurate panoramic depth with the assistance of other projection formats (e.g., Cubemap projection (CP) and Tangent projection (TP)). These projections significantly reduce distortions by mapping a panorama onto several tangent faces, but damage the inherent continuous representation of panoramic images. Many approaches (Wang et al., 2020; Jiang et al., 2021; Li et al., 2022; Ai et al., 2023) typically leverage the complementarity of these projection representations (shown in Fig. 1a) to estimate more accurate panoramic depth. Nevertheless, the trade-off between the FoV continuity and distortion elimination poses another challenge for the model design. Besides, these bi-projection fusion strategies require additional branches, increasing the computational burden. In view of this, we only take the equirectangular projection as the input and delve into the properties of the spherical space to deal with the distortion.
On the other hand, some approaches (Pintore et al., 2021; Shen et al., 2022a; Yun et al., 2023) deal with the panoramic distortions via modeling long-range dependencies (as illustrated in Fig. 1b). In previous work (Pintore et al., 2021), long short-term memory (Hochreiter and Schmidhuber, 1997) (LSTM) is introduced to disseminate global information in vertically compressed sequences. However, the geometric structure, which is an essential cue to estimate depth, is neglected, thus delivering blurred depth details. Meanwhile, those methods based on modeling long-range dependencies are either too global to perceive the local details (Yun et al., 2023), or too local to deal with the distortion efficiently (Shen et al., 2022a).
In this paper, we incorporate spherical geometry into our model (i.e., SGFormer, as shown in Fig. 1c) to capture local details while promoting global interaction. We leverage three geometric priors of the sphere to design a customized panoramic decoder (SPDecoder). Firstly, we use the spherical equidistortion to design a bipolar re-projection paradigm to deal with the distortions. We re-project these areas with severe distortions to the equator region to eliminate the negative effect of distortions. Especially, our method is parameter-free and transformation-efficient without any other branches. Secondly, we perform circular rotation through the continuity of the sphere. It helps to produce seamless panoramic representations, which allows two separate points, originally located on the left and right boundaries, to be close to each other and have similar semantics. Thirdly, with the spherical surface distance, curve local embedding is presented to generate curved relative position embedding for each point in the spherical projection space, which provides additional geometric distance and enhances the understanding of spatial structure.
Moreover, to further enhance performance comprehensively, we design a query-based global conditional position embedding (GCPE) scheme. Using features of different resolutions as queries to query the keys that match the resolution, we can adaptively provide different geometric details and spatial structures at different resolutions. This additional structure, in turn, explicitly guides our SPDecoder with global spatial positioning, thereby improving the accuracy of estimated depth.
Finally, we evaluate the proposed SGFormer on two benchmarks, Structured3D (Zheng et al., 2020) and Pano3D (Albanis et al., 2021). Experimental results demonstrate that our method outperforms existing methods by a large margin on both quantitative and qualitative results. Our main contributions can be summarized as follows:
  • We propose a spherical geometry transformer, named SGFormer, to integrate spherical structures and global positional perception simultaneously, contributing to superior performance over state-of-the-art (SoTA) solutions.
  • We design the SPDecoder by incorporating three spherical geometric priors: equidistortion, continuity, and surface distance, preserving the structural integrity of the sphere and enhancing the perception of spherical geometry and local details.
  • We introduce a query-based global conditional position embedding, which provides an explicit geometric cue, thereby sharpening the depth structure across various patches.
2. Related Work
2.1. Monocular $360^{\circ}$ Depth Estimation
Bi-projection fusion strategy. BiFuse (Wang et al., 2020) introduces a dual-branch fusion method that combines CP and ERP features to alleviate distortion, while UniFuse (Jiang et al., 2021) proposes a new framework that unidirectionally fuses CP features into ERP features during the decoding stage, enhancing both efficiency and performance. Glpanodepth (Bai et al., 2022) proposes a fusion module that combines ERP and CP features for panoramic depth estimation. 360MonoDepth (Rey-Area et al., 2022) introduces a framework that uses tangent images for depth estimation from high-resolution $360^{\circ}$ images. OmniFusion (Li et al., 2022) directly converts $360^{\circ}$ images into less-distorted perspective patches and uses geometric priors to merge the predictions of these patches into the ERP space to produce the final depth output in ERP format. HRDFuse (Ai et al., 2023) employs a collaborative mechanism that explores both the overall contextual information of the ERP and the local structure of the TP, estimating panoramic depth by learning the similarity of features between the ERP and TP.
Modeling long-range dependencies. SliceNet (Pintore et al., 2021) divides the ERP image into vertical slices, however, it is hard to recover the details near the poles because it ignores the latitudinal distortion property. PanoFormer (Shen et al., 2022a) attempts to estimate panoramic depth by leveraging a transformer-based architecture to process TP patches. By sampling the eight most relevant tokens for each central token on the ERP domain and adaptively adjusting local window sizes with a learnable offset, it enhances the precision in capturing the 2D structure and spatial details of each patch. EGFormer (Yun et al., 2023) leverages equirectangular geometry as the bias for local windows, thus estimating panoramic depth through an equirectangular geometry-aware local attention with a large receptive field. However, the vertical and horizontal shaped local windows also weaken the perception of the structure of each local area.
Other perspectives. Some researchers (Coors et al., 2018; Dai et al., 2017; Eigen et al., 2014; Zhu et al., 2020; Xiong and Grauman, 2018; Jiang et al., 2019; Khasanova and Frossard, 2019; Laina et al., 2016) employ various distortion-aware convolution filters to mitigate distortion, but due to the limited receptive fields, the final effect is still not entirely satisfactory. In deeper levels (Su and Grauman, 2019). Several spherical CNNs (Cohen et al., 2018; Yang et al., 2020) have employed convolution within the spectral domain, which could lead to a heavier computational burden. (Shen et al., 2022b) considers the significance of the geometric structure in depth estimation and proposes a neural contourlet network for panoramic monocular depth estimation.
2.2. Transformer for Vision Tasks
Figure 2. Overview of our proposed SGFormer. The SGFormer is composed of three major parts: feature extractors for ERP images at varying resolutions, calculation of global conditional position embedding (CGCPE) module, and the SPDecoder that integrates three spherical geometric priors.
Inspired by the transformer in the NLP field, the Vision Transformer (Dosovitskiy et al., 2020) (ViT) uses a global self-attention mechanism for visual tasks. However, the global information interaction often results in higher computational costs. Swin Transformer (Liu et al., 2021) (SwinT) adopts a window-based self-attention, incorporating a hierarchical architecture and shifting window mechanism for image processing. CSwin Transformer (Dong et al., 2022) (CSwinT) introduces a cross-shaped window attention mechanism, conducting parallel self-attention computations on horizontal and vertical stripes. (Ling et al., 2023; Cho et al., 2022) propose distortion relief methods specifically designed for panoramic image vision tasks. Despite these advancements, the distortion in panoramic images poses some challenges for transformers in panoramic image processing.
PanoFormer (Shen et al., 2022a) redesigns the traditional transformer with tangent patches on the spherical domain and panorama specific metrics, using learnable token flows to perceive geometric structures adaptively. HRDFuse (Ai et al., 2023) processes panoramic images into TP patches and aggregates global information in a ViT manner. EGFormer (Yun et al., 2023) combines with CSwinT (Dong et al., 2022)to propose an equirectangular geometry-biased transformer. However, this horizontal and vertical self-attention weakens the perception of local structures. In this work, we use the spherical geometric characteristics to further explore the suitable transformer structure for panoramic views.
3. SGFormer
3.1. Architecture Overview
Figure 3. Overview of the query-based CGCPE module. f3 generates a global key (i.e. GCPE3) by passing through a transformer module with global spherical position embedding (GSPE). The global key is then processed through linear layers to generate keys corresponding to features at various resolutions. Features of different resolutions are used as queries, which are subsequently operated by $QK^{T}$. After another linear transformation, the corresponding GCPEs at different resolutions are obtained.
In Fig. 2, we show our complete framework motivated by the spherical properties. Overall, the proposed SGFormer takes a panorama as input and outputs a corresponding high-quality depth map. During the encoder stage, we employ Resnet-34 to extract features at five different resolutions (Fig. 2 ①). Then these features are fed into the CGCPE module (Fig. 2 ② ), which is a query-based module that adaptively produces GCPEs, compensating for additional spatial geometric structure (More details are in Section 3.2). The visualization results of GCPEs demonstrate that they provide geometric details, promoting a comprehensive understanding of spatial location. Subsequently, a hierarchical decoder (the SPDecoder, described in Section 3.3 ) is designed to estimate depth from these extracted features and integrates three spherical geometric priors to maintain the unity of the spherical geometric structure. Due to explicit structural cues and spherical priors (i.e., equidistortion, continuity, and surface distance), our scheme can relieve distortion effectively and these structure- and sphere-aware designs can promote the performance of our model.
3.2. CGCPE Module
Previous work (Chen et al., 2021) has proved that the structure of the scenes is crucial for depth estimation. Hence, we aim to leverage the panoramic geometric cue to assist in the depth estimation process. To achieve this, we introduce GCPEs, which are learned from the extracted features and a spherical surface distance (i.e. Fig. 3 GSPE), to conduct the depth estimation with an additional explicit structure cue. The introduced GCPEs provide scale-adaptive variable details, which can push the network to explore the diverse spatial structures of the same panoramic space but with different resolutions.
Specifically, as shown in Fig. 3, we select the feature with a $\frac{H}{16}\times\frac{W}{16}$ resolution (denoted as $f_{3}$) as the reference feature. There is a trade-off between computational efficiency and clear details. To balance them, we choose the middle scale. The selected features are fed into a transformer module and then we achieve position embedding via a GSPE approach (more details are in section 3.3, curve local embedding), thereby promoting the generation of panoramic structure (i.e., global key in Fig. 3). As shown in the top-left corner of Fig. 3, GSPE projects any two pixels in the ERP feature map into the spherical space to calculate the spherical surface distance as a sphere-aware embedding. Subsequently, the aggregated global geometric information is transformed into keys via a linear projection layer. At each scale, we leverage features as queries to query the keys, adaptively generating GCPE at each resolution. From Fig. 4, the depth details can be observed clearly, indicating that the query-based GCPE strategy can contribute to the spatial structure and depth details.
Figure 4. The visualization of heatmap of features and the predicted depth map with/without GCPE.
Figure 5. Overview of SPDecoder (left), which is combined with three spherical geometric priors, and SPAttention (right).
3.3. SPDecoder
To balance the global perception and the local details enhancement, we propose the SPDecoder by leveraging the spherical priors. As illustrated in Fig. 5, we send the input features to SPAttention firstly, where we add curve local embedding to provide additional sphere-aware geometric distance. Then, considering the sphere’s continuity, we perform circular rotation to enhance the feature interaction among patches (Fig. 5 ①). We restore the original state through reverse circular rotation (Fig. 5 ②). Afterward, a bipolar re-projection approach based on the spherical equidistortion property is designed to relieve the distortion (Fig. 5 ③). Moreover, circular rotation is operated again to interact with the original features in the vertical direction (Fig. 5 ④). Ultimately, we revert to the original state through reverse operations (Fig. 5 ⑤⑥). We describe these operations in detail as follows.
Figure 6. (a) A projection imaging model with the x-axis (North) as the positive direction. (b) A projection imaging model with the y-axis (North) as the positive direction.
Bipolar re-projection (BRP): Since ERP images exhibit significant distortion in the polar regions and slight distortion near the equator areas, BRP can provide a content-consistent but perspective-different panorama via a bipolar transformation operation. Given a feature map $I\in R^{C\times H\times W}$, we implement BRP through a back projection operation. Specifically, we generate a canvas $I^{\prime}\in R^{C\times H\times W}$ and then transform the canvas from the ERP domain to a unit 3D Cartesian coordinates (Fig. 6 b). $\forall P\left(u^{\prime},v^{\prime}\right)\in I^{\prime}$, the transformation process can be denoted as follows:
(1) $\left\{\begin{matrix}\lambda^{\prime}=\frac{v^{\prime}}{H}\times\pi\\ \varphi^{\prime}=\frac{u^{\prime}}{W}\times 2\pi\end{matrix}\right.$
where $(u^{\prime},v^{\prime})$ indicates the position of the point P in ERP domain; $\lambda^{\prime}\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$/$\varphi^{\prime}\in\left(-\pi,\pi\right)$ represents latitude/longitude, respectively. Subsequently, we can calculate the $\left(x,y,z\right)$ coordinates corresponding to $\left(\lambda^{\prime},\varphi^{\prime}\right)$:
(2) $\left\{\begin{matrix}x=\sin\left(\lambda^{\prime}\right)\\ y=\cos\left(\lambda^{\prime}\right)\\ z=\cos\left(\frac{\pi}{2}-\lambda^{\prime}\right)\cdot\sin\left(\varphi^{\prime}\right)\end{matrix}\right.$
Similarly, the original feature map points can also be transformed into the same unit 3D Cartesian coordinates. And they have the following mapping relationships:
(3) $\left\{\begin{matrix}\lambda=\arccos\left(\frac{x}{\sqrt{z^{2}+y^{2}}}\right)\\ \varphi=\arcsin\left(\frac{y}{\sqrt{z^{2}+y^{2}}}\right)\end{matrix}\right.$
where $\lambda\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$ and $\varphi\in\left(-\pi,\pi\right)$ are the spherical coordinates of the original feature map (Fig. 6 a). Then we can obtain the position of point $P\left(u^{\prime},v^{\prime}\right)$ in the original feature map through the following formula:
(4) $\left(u,v\right)=\left(\frac{\lambda}{\pi}\times H,\frac{\varphi}{2\pi}\times W\right)$
In particular, we employ the bilinear interpolation function to fill the missing pixels. After BRP operation, the polar regions are projected to the equator (the first column of Fig. 7), which allows the model to estimate accurate depths with these distortion-free features. Besides, the proposed BRP approach is parameter-free and efficient in transformation, and no additional branches are required.
Figure 7. The first row shows the division of patches before and after CR. The second row displays the division of patches after BRP followed by another CR. The blue and red lines represent the divisions before and after CR, respectively.
Circular rotation (CR): The traditional transformer (Vaswani et al., 2017) architecture primarily relies on a global self-attention mechanism, which is inefficient when dealing with high-resolution images. To address this, SwinT (Liu et al., 2021) performs self-attention among divided window-based patches and employs a shifting window operation to facilitate feature interaction between the adjacent patches. Due to the unique seamless property of panoramic images, we implement self-attention interactions between local windows by rotating half the window size in the horizontal direction (the first row in Fig. 7). Then we rotate back to the original state to align with the initial spatial representation. Since the north and south poles of panoramic images are not adjacent in space, we cannot directly apply CR in the vertical direction. However, benefiting from the BRP module, we can map the features from the vertical to the horizontal direction. Then, through CR, features that are originally in the vertical direction are interacted. Finally, we obtain a feature representation that combines horizontal and vertical directions.
Figure 8. The left image shows an ERP format panoramic image, with each square indicating a segmented patch. Red and blue squares mark patches from the top and bottom with high distortion, while yellow and green squares indicate middle area patches with less distortion. These patches are projected onto the spherical space on the right.
Curve local embedding (CLE): In the ERP domain, whether in the polar regions with significant distortion or near the equator with less distortion, the distances between two points within a patch are relatively close regardless of the degree of distortion (on the left side of Fig. 8). However, after projecting onto the spherical domain, the distances between these points vary significantly according to the degree of distortion (on the right side of Fig. 8). The network can perceive the panoramic distortion distribution from the mutative distance between the two relevant points.
Hence, we introduce CLE to calculate the distance between two pixels in spherical space. In the SPAttention, we employ CLE as a relative position embedding to model the distortion distribution pattern, thereby delivering an explicit spherical geometry to make the model perceive the distortion easily.
For each patch containing N points, we define $C$ as the CLE within each ${N\times N}$ patch. $\forall P_{1}\left(\lambda_{1},\varphi_{1}\right)$ and $P_{2}\left(\lambda_{2},\varphi_{2}\right)$ (i.e., the position of two points within a patch in spherical coordinates) to calculate the CLE. $c\left(P_{1},P_{2}\right)$ represents the distance between $P_{1}$ and $P_{2}$, where $\lambda\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$/$\varphi\in\left(-\pi,\pi\right)$ represents latitude/longitude, repectively. Through the Haversine formula (Sinnott, 1984), we can calculate $c\left(P_{1},P_{2}\right)$ as follows:
(5) $h=\sin^{2}\left(\frac{\lambda_{2}-\lambda_{1}}{2}\right)+\cos\left(\lambda_{1}\right)\cdot\cos\left(\lambda_{2}\right)\cdot\sin^{2}\left(\frac{\varphi_{2}-\varphi_{1}}{2}\right)$
(6) $c\left(P_{1},P_{2}\right)=2R\cdot\arctan\left(\sqrt{\frac{h}{1-h}}\right)$
where the radius (R) is set to 1. We employ the Eq. 5 and Eq. 6 to obtain the CLE within a patch.
Table 1. Quantitative comparison on the S3D (Structured3D) and Pano3D datasets with the SoTA methods. Blue indicates that our method achieves the best performance.
Dataset Method Abs.rel $\downarrow$ Sq.rel $\downarrow$ RMS.lin $\downarrow$ RMSlog $\downarrow$ $\delta^{1}$uparrow $\delta^{2}$uparrow $\delta^{3}$uparrow
Bifuse (Wang et al., 2020) 0.0644 0.0565 0.4099 0.1194 0.9673 0.9892 0.9948
SliceNet (Pintore et al., 2021) 0.1103 0.1273 0.6164 0.1811 0.9012 0.9705 0.9867
Yun et al. (Yun et al., 2022) 0.0505 0.0499 0.3475 0.1150 0.9700 0.9896 0.9947
S3D (Zheng et al., 2020) PanoFormer (Shen et al., 2022a) 0.0394 0.0346 0.2960 0.1004 0.9781 0.9918 0.9958
EGFormer (Yun et al., 2023) 0.0342 0.0279 0.2756 0.0932 0.9810 0.9928 0.9964
SGFormer (ours) 0.0303 0.0293 0.2429 0.0863 0.9857 0.9940 0.9969
Bifuse (Wang et al., 2020) 0.1704 0.1528 0.7272 0.2466 0.7680 0.9251 0.9731
SliceNet (Pintore et al., 2021) 0.1254 0.1035 0.5761 0.1898 0.8575 0.9640 0.9867
Yun et al. (Yun et al., 2022) 0.0907 0.0658 0.4701 0.1502 0.9131 0.9792 0.9924
Pano3D (Albanis et al., 2021) PanoFormer (Shen et al., 2022a) 0.0699 0.0494 0.4046 0.1282 0.9436 0.9847 0.9939
EGFormer (Yun et al., 2023) 0.0660 0.0428 0.3874 0.1194 0.9503 0.9877 0.9952
SGFormer (ours) 0.0583 0.0364 0.3537 0.1101 0.9613 0.9896 0.9957
3.4. Loss Function
For a fair comparison, we strictly follow EGFormer (Yun et al., 2023), using the scale-and-shift-invariant loss function (Ranftl et al., 2020). Specifically, we align the predicted depth $D_{p}$ and the ground truth depth $D_{g}$ with their scale (s) and shift (t) appropriately through a least-squares criterion:
(7) $s,t=\arg\min\limits_{s,t}\left(s\cdot D_{p}+t-D_{g}\right)$
[…]
where $D_{p}^{\prime}$ and $D_{g}^{\prime}$ are the aligned predicted depth and the corresponding ground truth. Subsequently, we calculate the mean squared error loss ($L_{pix}$) as follows:
(9) $m=\left(D_{g}^{\prime}>0\right)$
(10) $L_{pix}=\frac{1}{n}\sum_{i=1}^{n}\left[m\cdot\left|\left(D_{p}^{\prime}-D_{g}^{\prime}\right)\right|\right]$
where the mask (m) is employed to filter out the invalid values in ground truth. Besides, the gradient loss ($L_{grad}$) is employed to sharpen the depth edges, which is calculated as follows:
(11) $L_{grad}=\sum_{x,y}\left(\left|G_{x}\left(D_{p}^{\prime}-D_{g}^{\prime}\right)_{x,y}\right|+\left|G_{y}\left(D_{p}^{\prime}-D_{g}^{\prime}\right)_{x,y}\right|\right)$
where $G_{x}$ and $G_{y}$ are the horizontal/vertical gradients of a predicted depth map. Ultimately, the total objective function can be written as:
(12) $L_{total}=L_{pix}+0.5\cdot L_{grad}$
Figure 9. Qualitative results on the Structured3D dataset (the top two lines), and Pano3D dataset (the bottom two lines). More results can be found in Supplementary Materials.
4. Experiments
We strictly follow the training and evaluation protocol of EGformer (Yun et al., 2023) to evaluate our proposed SGFormer approach with two solid benchmarks. The implementation details are described as follows.
4.1. Dataset and Related Details
The two datasets employed for evaluation: Structured3D (Zheng et al., 2020) and Pano3D (Albanis et al., 2021). Structured3D is a high-quality synthetic dataset designed specifically to support a variety of structured 3D modeling tasks, providing many realistic images with detailed 3D structure annotations. On the other hand, the Pano3D dataset serves as a new benchmark for spherical panorama depth estimation. It aims to comprehensively evaluate the performance of depth prediction models, including the accuracy, the clarity of edges, and the smoothness of the image.
The reasons for not using other datasets are as follows: Firstly, the PanoSunCG dataset (Wang et al., 2018) is no longer publicly available. Secondly, the images in the 3D60 dataset (Zioulis et al., 2018) exhibit depth value leakage through their pixel brightness. Thirdly, Pano3D incorporates a rendered version of the Matterport3D (Chang et al., 2017), employing Pano3D and Matterport3D for assessment would result in unnecessary duplication. Lastly, in the Stanford2D3D dataset (Armeni et al., 2017), the top and bottom parts of equirectangular images are not rendered properly. However, our method requires converting the north and south poles to the equatorial region through BRP. More details are included in the Supplementary Materials.
In our experimental settings, we perform all experiments using a single GTX 3090 GPU, where the batch size is set to 1. We choose the AdamW (Loshchilov and Hutter, 2017) as the optimizer. On the Structured3D dataset, the initial learning rate is set at $5\cdot 10^{-5}$, with the model being trained for 50 epochs. Subsequently, after merging the Pano3D and Structured3D datasets, the model continues to be further trained for 20 epochs., maintaining the same learning rate, and adopting a learning rate exponential decay strategy of 0.95 to optimize the training effect.
4.2. Comparison Results
Metrics: To evaluate the model performance fairly, we select a series of standard evaluation metrics in EGFormer (Yun et al., 2023), including absolute relative error (Abs.rel), squared relative error (Sq.rel), root mean square linear error (RMS.lin), and root mean square log error (RMSlog). The lower the value of these indicators, the better the performance of the model. Additionally, we use relative accuracy ($\delta^{n}$) as a measure of performance, where n represents different threshold levels (i.e., $\delta<1.25^{n}$). The higher the value of this indicator, the higher the accuracy of the predictions.
Quantitative Analysis: To validate the effectiveness of our approach, we compare our SGFormer model with the current state-of-the-art methods, including strategies for bi-projection fusion, and long-range dependencies. As shown in Table 1, our model ranks first in nearly all evaluation metrics, indicating superior overall performance. Specifically, on the one hand, compared with Bifuse (Wang et al., 2020) based on bi-projection fusion, our model has achieved a 53% improvement (Abs.rel) and 41% (RMS.lin) on the Structured3D dataset. For the Pano3D dataset, the performance is significantly improved by 66% (Abs.rel) and 51% (RMS.lin). On the other hand, compared with EGFormer (Yun et al., 2023) based on long-range dependencies, our model achieves an 11% improvement (Abs.rel) and 12% (RMS.lin) on the Structured3D dataset. For the Pano3D dataset, SGFormer achieves 12% (Abs.rel) and 9% (RMS.lin) gain. In quantitative comparison, we achieve significant performance improvements because we fully leverage the geometric priors of the sphere to relieve the panoramic distortion. The lower quality of depth estimation in SliceNet (Pintore et al., 2021) is mainly due to the detail loss when processing equirectangular images in the vertical direction.
Qualitative Analysis: In Fig. 9, we present a qualitative comparison with EGFormer (Yun et al., 2023) and PanoFormer (Shen et al., 2022a). We can observe that our performance is superior in terms of regional structural details. Specifically, as demonstrated in the red boxes in Fig. 9, more accurate depths are recovered. We perform better on local details because we provide additional spherical geometric structures.
Table 2. Ablation studies of individual components. Starting from a baseline approach without any improvements, we add components one at a time. We conduct a series of ablation studies on the S3D (Structured3D) and Pano3D datasets. Results in bold indicate the best results. (CLE: curve local embedding, GCPE: global conditional position embedding, CR: circular rotation, BRP: bipolar re-projection)
Data ID CLE GCPE CR BRP Abs.rel $\downarrow$ Sq.rel $\downarrow$ RMS.lin $\downarrow$ RMSlog $\downarrow$ $\delta^{1}$uparrow
a) 0.0388 0.0352 0.2838 0.0968 0.9800
b) 0.0362 0.0333 0.2652 0.0924 0.9822
S3D (Zheng et al., 2020) c) 0.0329 0.0307 0.2561 0.0898 0.9841
d) 0.0301 0.0299 0.2499 0.0874 0.9849
e) 0.0303 0.0293 0.2429 0.0863 0.9857
a) 0.0720 0.0483 0.4048 0.1278 0.9417
b) 0.0708 0.0436 0.3814 0.1238 0.9474
Pano3D (Albanis et al., 2021) c) 0.0649 0.0416 0.3774 0.1183 0.9543
d) 0.0626 0.0382 0.3619 0.1117 0.9598
e) 0.0583 0.0364 0.3537 0.1101 0.9613
Figure 10. Qualitative comparisons about individual components. The top row shows visual comparisons of depth estimation after adding different components, while the bottom row shows the visual comparisons of the corresponding error maps for the estimated depth maps. The middle two rows display close-up views of the highlighted areas in both the top and bottom rows respectively. As we add more components from left to right, the object boundaries become clearer, and the overall depth estimation becomes more accurate with lower errors. The trend of error changes can be directly observed from the error maps. a, b, c, d, e are the same as Table 2.
4.3. Ablation study
We conduct a thorough validation of the key components of the SGFormer model under the same experimental conditions. As shown in Table 2, initially, we establish a baseline with a model that does not include any specific components, using a Resnet-34 as the encoder and a ViT-based (Dosovitskiy et al., 2020) decoder. Then we sequentially introduce core components on the Structured3D and Pano3D datasets to explore the effectiveness of our model.
Effectiveness of CLE: From Table 2, we can observe that on the Structured3D dataset, there is a 6.70% improvement (Abs.rel) and 6.55% (RMS.lin). On the Pano3D dataset, the performance is further boosted by 1.67% (Abs Rel) and 5.78% (RMS.lin). This illustrates that by incorporating CLE, we introduce additional spherical surface distance, which aids the network in perceiving the distribution of distortion.
Effectiveness of GCPE: The introduction of GCPE compensates for the model’s geometric details and spatial structure across different resolutions. Results in Table 2 show that the performance is significantly improved by 9.12% (Abs.rel) and 3.43% (RMS.lin) on the structured3D dataset. For the Pano3D dataset, there is an improvement of 8.33% (Abs.rel) and 1.05% (RMS.lin).
Effectiveness of CR: The CR component is added to achieve cross-window feature interaction. Comparisons in Table 2 reveal an 8.51% improvement (Abs.rel) and 2.42% (RMS.lin) on the Structured3D dataset. The performance is further enhanced by 3.54% (Abs.rel) and 4.11% (RMS.lin) on the Pano3D dataset.
Effectiveness of BRP: The BRP component is incorporated to reduce distortion by projecting the north and south polar regions near the equator. Results from Table 2 show that performance is improved by 2.80% (RMS.lin) on the Structured3D dataset. For the Pano3D dataset, the performance is further boosted by 6.87% (Abs.rel) and 2.27% (RMS.lin).
Fig. 10 shows the results of a qualitative comparison. As we gradually add CLE, GCPE, CR, and BRP components from left to right, the output depth maps not only perform better in overall depth estimation but also capture local structural details more precisely. Furthermore, through the visualization comparison of error maps, we can clearly see that the estimated error is continuously decreasing (The deeper the red color, the higher the error. The deeper the blue color, the lower the error).
5. Conclusion
In this paper, to remove the adverse impacts of distortion, we integrate three spherical geometric priors: equidistortion, continuity, and surface distance, to propose a spherical geometry transformer, named SGFormer. Specifically, we employ BRP to project the polar regions with severe distortions towards the equator regions to relieve the distortion. This strategy is parameter-free and efficient without additional branches and complex transformation costs. CR leverages the continuity of the sphere, facilitating cross-window feature interaction and enhancing the understanding of the global structure. CLE generates relative position embeddings based on the pixel points projected onto the spherical space, providing additional spherical surface distance to perceive distortion explicitly. Additionally, we introduce a query-based global conditional position embedding strategy, which can dynamically provide varying spatial structure and geometric local details across different resolutions. Experiments on two popular benchmarks demonstrate that our proposed SGFormer significantly outperforms the SoTA methods.